Two samples are taken with the following numbers of successes and sample sizes r1 = 34 r2 = 24 n1 = 75 n2 = 63 Find a 93% confidence interval, round answers to the nearest thousandth.
Answer #1

Here, , n1 = 75 , n2 = 63

p1cap = X1/N1 = 34/75 = 0.4533
p1cap = X2/N2 = 24/63 = 0.381

p1cap = 0.4533 , p2cap = 0.381


Standard Error, sigma(p1cap - p2cap),
SE = sqrt(p1cap * (1-p1cap)/n1 + p2cap * (1-p2cap)/n2)
SE = sqrt(0.4533 * (1-0.4533)/75 + 0.381*(1-0.381)/63)
SE = 0.084

For 0.93 CI, z-value = 1.81
Confidence Interval,
CI = (p1cap - p2cap - z*SE, p1cap - p2cap + z*SE)
CI = (0.4533 - 0.381 - 1.81*0.084, 0.4533 - 0.381 + 1.81*0.084)
CI = (-0.0797 , 0.2243)